# JWST/NIRSpec insights into the circumnuclear region of Arp 220: A detailed kinematic study

### v1

@@ 5.3 Outflows properties and energetics
In this section, we estimate the outflow energetics of the molecular and of the ionised gas. To estimate the mass of the outflowing gas in the hot molecular phase, we followed Scoville et al. (1982) and Mazzalay et al. (2013):
@@
To compute the ionised outflow mass, we adopted the simplified model by Genzel et al. (2011), valid for case B recombination of fully ionized gas with K. The mass of outflowing gas is given by:
@@
The determination of the ionized gas mass is significantly influenced by , which can span a few orders of magnitude, from 102 to 104cm-3 (e.g. Villar Martín et al. 2014; Perna et al. 2017; Förster Schreiber et al. 2019; Isobe et al. 2023). We derived the electron density through the line ratio [Fe ii] 1.677 m/[Fe ii] 1.644 m. We fitted the integrated spectrum around the lines of interest considering in the model also the Br11 line, since it is blended with [Fe ii] 1.677 m. Since the latter line is very faint at spaxel level, we integrated the spectra over the individual outflow regions SEO (Fig. 6), HGO (Fig. 8), NWB (Fig. 10), and WNO (Fig. 12, bottom panels), to infer average electron densities for each outflow. Moreover, due to the low S/N of [Fe ii] 1.677, we considered the total profile of [Fe ii] 1.677 m and 1.644 m, without distinguishing the contribution of non-outflowing components. We obtained electron densities in the range log (ne/cm-3) from the relation in Koo et al. (2016).
We found that the region with the highest electron density is the WNO (n cm-3). For the HGO, we found n cm-3. We also integrated the spectrum and computed the electron density in both the red and blue-shifted region of the HG outflow, separately. We found n cm-3 for the red-shifted and a 3 upper limit of 5 cm-3 in the blue-shifted part. In the SEO and in the NWB the emission of [Fe ii] 1.677 m is not detected, resulting in upper limits on the electron density: cm-3, and cm-3, respectively. All the inferred values are reported in Table 3.
The estimates through the [Fe ii] transitions are higher (a factor 10-100) than those found from the [S ii] doublet in the optical band with MUSE (170 cm-3, Perna et al. 2020). This may indicate a gas stratification in the nuclear regions of Arp 220, with [Fe ii] emission, having higher critical density (Nc cm-3), tracing the densest part of the ISM, and [S ii] (Nc cm-3) tracing the more diffuse gas. We chose to estimate the outflow properties with both electron density estimations from [Fe ii] and [S ii].
@@
where C is a factor that depends on the outflow history. We adopted a constant mass outflow rate (C=1). We calculated the radius as the maximum extension from the corresponding nucleus. Specifically, for the HGO and WNO, we calculated the distances from the W nucleus, while for the SEO and NWB we calculated them from the E nucleus. We associated the NWB to the E nucleus, in order to have a more conservative estimation of the mass outflow rate. This choice is also motivated by the fact that the outflows launched by the W nucleus are oriented along the NS direction (Sect. 4.2), while the one of the E nucleus is along the SE direction. Moreover, the 100 pc scale molecular outflow launched by the E nucleus and traced by CO (Wheeler et al., 2020) is similarly oriented along the SE-NW direction.
@@
where and are the velocities relative to the systemic velocity of the nucleus from which the outflow originates, and the velocity dispersion of the outflowing component, respectively. We computed the outflow velocities as the flux-weighted average of the outflow velocity of each spaxel. The properties of the ionised and molecular gas outflows for the analyzed regions are reported in Table 3.
The outflow masses and the mass outflow rates of the hot molecular gas vary from 700 M⊙ in the SEO to 2000 M⊙ in the HGO, and from M⊙ yr-1 for the NWB to M⊙ yr-1 for the HGO, respectively. For the ionized gas, the outflow mass varies from a minimum of M⊙ for the SE outflow to a maximum of M⊙ fort the WNO, calculated with the density from the [Fe ii] doublet. The highest mass outflow rate of ionised gas is measured in the WNO with a value of 8 M⊙ yr-1, estimated using the density from the [S ii], while the lowest is measured in the NWB with a value of 0.4 M⊙ yr-1.
@@ 6 Discussion
Cold molecular outflows have been detected in the nuclear region of Arp 220 with different emission line tracers, such as HCO+, H2O, H2O+, OH, and OH+ (Sakamoto et al. 2009; González-Alfonso et al. 2012; Veilleux et al. 2013; Tunnard et al. 2015; Martín et al. 2016; Zschaechner et al. 2016). However, most of them are unresolved and therefore their morphology could not be constrained. Spatially resolved outflows studies have been conducted by Barcos-Muñoz et al. (2018a) and Wheeler et al. (2020) with ALMA, and Perna et al. (2020) with MUSE. Thanks to the possibility of resolving multiphase outflows with NIRSpec, we can compare the outflow properties of hot molecular and ionized phases derived in this work, with those obtained from previous observations for cold molecular gas and ionized gas at larger scale. This will help us to understand the connection between multiphase and multiscale outflows. After discussing the structures of these outflows, we will examine the mechanisms responsible for launching them.
6.1 W nucleus driven outflows
@@
Barcos-Muñoz et al. (2018b) reported a cold molecular outflow launched by the W nucleus, traced by HCN(1-0), extending for 120 pc, with a deprojected velocity of 850 km s-1 and a lower limit of the H2 mass for the red-shifted and blue-shifted component of and , respectively. These lower limits consider the possible presence of additional outflowing material at lower velocities, not accounted due to its overlap with the disc emission. Wheeler et al. (2020) found an upper limit for the total H2 molecular mass of for the redshifted outflow, and for the blue-shifted, traced by CO(3–2) gas. The upper limit here accounts the possible contamination of H13CN (4–3) emission, very close to the CO(3–2) transition.
@@
We detected an HG-shape outflow originating from the W nucleus extending for 600 pc, with a hot H2 outflowing gas mass of 2000 M⊙. The redshifted region accounts for 1200 M⊙, while the blue-shifted for 800 M⊙. This outflow is in the same direction as the cold molecular outflow detected in Wheeler et al. (2020), with receding gas northward and approaching gas southward. Therefore, we can reasonably assume that cold and hot molecular outflows are connected and associated with the same outflow event. Under this assumption, the total mass of the hot and cold molecular gas result in an estimated hot-to-cold molecular gas ratio of , obtained considering both receding and approaching gas. This order-of-magnitude estimate is consistent with the values found in nearby ULIRGs (Pereira-Santaella et al., 2016; Emonts et al., 2017; Pereira-Santaella et al., 2018; Lamperti et al., 2022).
@@
Comparing the total outflowing mass of ionized and total molecular gas in the W nucleus, hence accounting for both HGO and WNO, we report a mass outflow ratio between the two phases of 0.16 assuming [S ii]-based electron densities ( with the density estimated with [Fe ii]). This is not unusual since the molecular phase usually dominates the mass outflow rate budget, exceeding the ionised gas phase by 1-3 order of magnitude (Rupke & Veilleux 2013; García-Burillo et al. 2015; Carniani et al. 2015; Fiore et al. 2017; Cresci et al. 2023).
@@
Perna et al. (2020) identified an extended structure associated with high-velocity ionized gas (see their Fig. 19, velocity channel 125 km s-1 ¡ —v— ¡ 375km s-1) from northeast to southwest, that appears to be aligned with the hourglass-shaped outflow launched from the W nucleus. This suggests that the HGO could even be connected to the outflow detected at larger scale. Alternatively, they could trace different ejection events along the same direction.
6.2 E nucleus driven outflows
@@
Wheeler et al. (2020) reported a collimated and extended molecular outflow in the E nucleus traced by CO transitions (J). The outflow found in the present work (SEO), launched from the E nucleus, is consistent with the direction of the outflow observed in the cold molecular gas, blue-shifted towards the south direction. Wheeler et al. (2020) reported a mass for the blueshifted outflow of 2.7 M⊙, that, under the assumption that the CO outflow is associated with the SE outflow detected in our data would lead to a hot-to-cold molecular gas ratio of , compatible with the value we inferred in Sect. 6.1 for the outflow launched by the W nucleus.
@@
We attributed the NW bubble to the E nucleus, since this structure could represent the counterpart of the blushifted outflow detected in the SE direction. However, unlike the HGO, in this case there is no clear blushifted and redshifted part of the outflow. This may be due to different gas ejection events, or to the fact that the redshifted part of the outflow is more obscured. Another possibility is that we included part of the red-shifted outflow launched from the E nucleus in the HGO, since the two outflows may overlap, as we can see in Fig. 5c.
@@
We find that the mass of the ionised gas in the NW bubble, assuming ne = 170 cm-3, is M⊙, with a total energy of erg. Lockhart et al. (2015) estimated the properties of the NW bubble using H+[NII] narrow band imaging obtained with HST/WC3 with a resolution of 0.0396”/pixel, corresponding to 15 pc. They used assumptions different from ours on the density, on velocity of the bubble and also on the geometry. In fact, they adopted ne = 10 cm-3, about one order of magnitude lower than that used in this work, a velocity of 250 km s-1, around a factor 2 lower and a shell-like geometry. Despite these different assumptions, their derived outflow mass ( M⊙) and energy ( erg) are in agreement, within a factor of two, with our results.
@@
We also compared the outflows detected with NIRSpec with the larger scale outflows observed with MUSE in Perna et al. (2020). In that case, the most extreme velocities for the ionised outflow are observed along PA = 138° 222Measured anti-clockwise from the north.; these extreme motions are aligned with X-ray emission and with the minor axis of the stellar disc. The SEO and the NWB are found along the same PA, but at ¡ 1 kpc scales, that might suggest that they are related to the kpc-scale atomic outflow presented in Perna et al. (2020). The ionized, large-scale bubble observed in MUSE (kpc) and the NW bubble could be, also in this case, the results of different outflow episodes. These multiple bubbles at different distances are very similar to those predicted by the TNG50 cosmological simulation as a result of AGN feedback (Nelson et al. 2019) and to those observed in NGC 1275 (e.g. Fabian et al. 2003; Sanders & Fabian 2007) and in the Tea Cup (Venturi et al. 2023).
6.3 What drives the outflows?

### v2

@@ 5.3 Outflows properties and energetics
In this section, we estimate the outflow energetics of the molecular and of the ionized gas. To estimate the mass of the outflowing gas in the hot molecular phase, we followed Scoville et al. (1982) and Mazzalay et al. (2013):
@@
To compute the ionized outflow mass, we adopted the simplified model by Genzel et al. (2011), valid for case B recombination of fully ionized gas with K. The mass of outflowing gas is given by
@@
The determination of the ionized gas mass is significantly influenced by , which can span a few orders of magnitude, from 102 to 104cm-3 (e.g., Villar Martín et al. 2014; Perna et al. 2017; Förster Schreiber et al. 2019; Isobe et al. 2023). We derived the electron density through the line ratio [Fe ii] 1.677 m/[Fe ii] 1.644 m. We fitted the integrated spectrum around the lines of interest considering in the model also the Br11 line, since it is blended with [Fe ii] 1.677 m. Since the latter line is very faint at spaxel level, we integrated the spectra over the individual outflow regions SEO (Fig. 6), HGO (Fig. 8), NWB (Fig. 10), and WNO (Fig. 12, bottom panels), to infer average electron densities for each outflow. Moreover, due to the low S/N of [Fe ii] 1.677, we considered the total profile of [Fe ii] 1.677 m and 1.644 m, without distinguishing the contribution of non-outflowing components. We obtained electron densities in the range log (ne/cm-3) from the relation in Koo et al. (2016).
We found that the region with the highest electron density is the WNO (n cm-3). For the HGO, we found n cm-3. We also integrated the spectrum and computed the electron density in both the red- and blueshifted region of the HG outflow, separately. We found n cm-3 for the redshifted and a 3 upper limit of 5 cm-3 in the blueshifted part. In the SEO and in the NWB the emission of [Fe ii] 1.677 m is not detected, resulting in upper limits on the electron density: cm-3, and cm-3, respectively. All the inferred values are reported in Table 3.
The estimates through the [Fe ii] transitions are higher (a factor 10-100) than those found from the [S ii] doublet in the optical band with MUSE (170 cm-3, Perna et al. 2020). This may indicate a gas stratification in the nuclear regions of Arp 220, with [Fe ii] emission, having higher critical density (Nc cm-3), tracing the densest part of the ISM, and [S ii] (Nc cm-3) tracing the more diffuse gas. It is worth noting that using a constant density of 102 cm-3, as the one derived by Perna et al. (2020) using [S ii], could lead to an overestimation of the gas masses and the energetics, because this line emission likely traces partially ionized gas with lower densities (Davis et al. 2011; Revalski et al. 2022). On the other side, our estimation of the from [Fe ii] doublet comes only from the nuclear (WNO) and the most compressed regions (HGO), where the faint [Fe ii] 1.677 m line is detected. Thus, this electron density may not be representative of the density in the full FoV. Therefore, we chose to estimate the outflow properties with electron density estimations from both [Fe ii] and [S ii] that may represent reasonable lower and upper boundaries.
@@
where C is a factor that depends on the outflow history. We adopted a constant mass outflow rate (C=1). We calculated the radius from the flux maps of each outflow. In particular, we measured as the maximum extension from the corresponding nucleus. Specifically, for the HGO and WNO, we calculated the distances from the W nucleus, while for the SEO and NWB we calculated them from the E nucleus. We associated the NWB to the E nucleus, in order to have a more conservative estimation of the mass outflow rate. This choice is also motivated by the fact that the outflows launched by the W nucleus are oriented along the NS direction (Sect. 4.2), while the one of the E nucleus is along the SE direction. Moreover, the 100 pc scale molecular outflow launched by the E nucleus and traced by CO (Wheeler et al., 2020) is similarly oriented along the SE-NW direction.
@@
where and are the velocities relative to the systemic velocity of the nucleus from which the outflow originates, and the velocity dispersion of the outflowing component, respectively. We computed the outflow velocities as the flux-weighted average of the outflow velocity of each spaxel. We assigned the standard deviation of the values in that region as the uncertainty on vout. The properties of the ionized and molecular gas outflows for the analyzed regions are reported in Table 3.
The outflow masses and the mass outflow rates of the hot molecular gas vary from 700 M⊙ in the SEO to 2000 M⊙ in the HGO, and from M⊙ yr-1 for the NWB to M⊙ yr-1 for the HGO, respectively. For the ionized gas, the outflow mass varies from a minimum of M⊙ for the SE outflow to a maximum of M⊙ fort the WNO, calculated with the density from the [Fe ii] doublet. The highest mass outflow rate of ionized gas is measured in the WNO with a value of 8 M⊙ yr-1, estimated using the density from the [S ii], while the lowest is measured in the NWB with a value of 0.4 M⊙ yr-1.
@@
5.4 Three-dimensional outflow modeling
To measure the geometrical parameters of SEO, HGO and the NWB outflows we used the modeling framework called Modeling Outflows and Kinematics of AGN in 3D () presented in Marconcini et al. (2023). By assuming a geometry and a simple constant radial velocity field, the model is able to retrieve the outflow inclination with respect to the line of sight (LOS), and the intrinsic outflow velocity. Full details and the implementations of the model are discussed in Marconcini et al. (2023). Here we just show the results of the model for the hot molecular gas, by using a conical geometry for the SEO and two expanding bubbles for HGO and NWB, assuming a constant radial velocity (See Appendix B). We also assumed an intrinsic dispersion velocity of 30 km s-1 to account for the turbulent nature of the bubbles. Since we found that the ionized gas has a similar kinematics, the results of the model for the ionized phase would be the same. We did not attempt to model the WNO since it is very compact and we could not establish a clear geometry to be assumed.
In Fig. 19 we show the 3D model of the outflows detected in the nuclear region of Arp 220. The geometry parameters derived from are as follows.
The aperture of the SEO is 80° with an inclination respect to the LOS of 70° and PA of 230° measured clockwise from north. This is consistent with a conical outflow launched perpendicular from the E circumnuclear disk having an almost edge-on orientation, according to Wheeler et al. (2020). The velocity retrieved for the SEO by is 700 km s-1, consistent within with the velocities computed as defined as in Eq. 6), but slightly higher.
The HGO has an inclination of –60° with respect to the LOS and a velocity of 400 km s-1 for the blueshifted side, and 120° and 600 km s-1 for the redshifted side. The discrepancies between the velocity in the redshifted and blueshifted side could be explained by the possible ISM porosity and hence different propagation of the bubbles. The inclination of the W circumnuclear disk is 120° with respect to the LOS (Wheeler et al. 2020), which mean that the axis of the bubble is not exactly perpendicular to the disk, but has an offset of about 30°.
For the NWB we found an inclination of 68° with respect to the LOS and a velocity of 750 km s-1, also in this case slightly higher than our flux-weighted velocity measurement.
In all outflows, the velocities defined by Eq. 6 seem to underestimate the velocities found by the even though they are consistent within 1-2. This is due to the fact that the model takes into account the intrinsic velocity of the outflow corrected for all projection effects. In conclusion, the kinematic outflow properties inferred with and the method outlined in Sect.5.3 are consistent within the uncertainties.
@@ 6 Discussion
The ubiquity of multiphase outflows in ULIRGs was established through spatially resolved observations in the visible band, thanks to integral field instruments like VIMOS and MUSE (Arribas et al. 2014; Westmoquette et al. 2012; Venturi et al. 2021; Musiimenta et al. 2024), in the near IR band with Gemini/NIFS, Keck/OSIRIS, and VLT/SINFONI (Engel et al. 2011a; Riffel et al. 2015; Storchi-Bergmann et al. 2010; Pereira-Santaella et al. 2016; Emonts et al. 2017; U et al. 2019; Schönell et al. 2019), and in the radio band with ALMA (Pereira-Santaella et al. 2016).
Simulations show that outflows likely contain cold, dense clouds that are accelerated by a stream of warm and hot gas (e.g., Wagner et al. 2012; Zubovas & King 2014). However, the direct comparison of outflows identified with different tracers (and facilities), required to test such predictions, has been so far limited to very few targets. A key challenge in studying the direct impact of outflows on the ISM is the need for both high sensitivity and spatial resolution. Ground-based near-IR integral field facilities like SINFONI, OSIRIS and NIFS mainly traced the hot molecular (e.g., H2) and ionized (Br) gas with high spatial and spectral resolution. However, due to limitations in sensitivity and the presence of airglow emission lines, it has not always possible to fully characterize these outflows and compute directly their properties (e.g Riffel et al. 2014; Emonts et al. 2017; U et al. 2019; Bianchin et al. 2022). This is critical in order to characterize the connection between the warm molecular and the ionized gas outflow with the cold molecular gas. In fact, thanks to the advent of millimeter/submillimeter telescopes such as ALMA, high-spatial-resolution images of cold molecular gas in nearby galaxies have enabled us to reveal the structure and the morphology of starburst-driven and AGN-driven molecular outflows in high details (e.g Salak et al. 2020; Lamperti et al. 2022; Holden et al. 2024).
6.1 The multiphase outflows
@@
Cold molecular outflows have been detected in the nuclear region of Arp 220 with different emission line tracers, such as HCO+, H2O, H2O+, OH, and OH+ (Sakamoto et al. 2009; González-Alfonso et al. 2012; Veilleux et al. 2013; Tunnard et al. 2015; Martín et al. 2016; Zschaechner et al. 2016). However, most of them are unresolved and therefore their morphology could not be constrained. Spatially resolved outflows studies have been conducted by Barcos-Muñoz et al. (2018a) and Wheeler et al. (2020) with ALMA, and Perna et al. (2020) with MUSE. Thanks to the possibility of resolving multiphase outflows with NIRSpec, we can compare the outflow properties of hot molecular and ionized phases derived in this work, with those obtained from previous observations for cold molecular gas and ionized gas at larger scale. This will help us understand the connection between multiphase and multiscale outflows. After discussing the structures of these outflows, we examine the mechanisms responsible for launching them.
6.1.1 W nucleus-driven outflows
@@
Barcos-Muñoz et al. (2018b) reported a cold molecular outflow launched by the W nucleus, traced by HCN(1-0), extending for 120 pc, with a deprojected velocity of 850 km s-1 and a lower limit of the H2 mass for the redshifted and blueshifted component of and , respectively. These lower limits consider the possible presence of additional outflowing material at lower velocities, not accounted due to its overlap with the disk emission. Wheeler et al. (2020) found an upper limit for the total H2 molecular mass of for the redshifted outflow, and for the blueshifted, traced by CO(3–2) gas. The upper limit here accounts the possible contamination of H13CN (4–3) emission, very close to the CO(3–2) transition.
@@
We detected an HG-shape outflow originating from the W nucleus extending for 600 pc, with a hot H2 outflowing gas mass of 2000 M⊙. The redshifted region accounts for 1200 M⊙, while the blueshifted for 800 M⊙. This outflow is in the same direction as the cold molecular outflow detected in Wheeler et al. (2020), with receding gas northward and approaching gas southward. Therefore, we can reasonably assume that cold and hot molecular outflows are connected and associated with the same outflow event. Under this assumption, the total mass of the hot and cold molecular gas result in an estimated hot-to-cold molecular gas ratio of , obtained considering both receding and approaching gas. This order-of-magnitude estimate is consistent with the values found in nearby ULIRGs (Pereira-Santaella et al., 2016; Emonts et al., 2017; Pereira-Santaella et al., 2018; Lamperti et al., 2022).
@@
After comparing the total outflowing mass of ionized and total molecular gas in the W nucleus and hence accounting for both HGO and WNO, we report a mass outflow ratio between the two phases of 0.16 assuming [S ii]-based electron densities ( with the density estimated with [Fe ii]). This is not unusual since the molecular phase usually dominates the mass outflow rate budget, exceeding the ionized gas phase by one to three orders of magnitude (Rupke & Veilleux 2013; García-Burillo et al. 2015; Carniani et al. 2015; Fiore et al. 2017; Cresci et al. 2023).
@@
Perna et al. (2020) identified an extended structure associated with high-velocity ionized gas (see their Fig. 19, velocity channel 125 km s-1 ¡ —v— ¡ 375km s-1) from northeast to southwest, that appears to be aligned with the hourglass-shaped outflow launched from the W nucleus. This suggests that the HGO could even be connected to the outflow detected at larger scales. Alternatively, they could trace different ejection events along the same direction.
6.1.2 E nucleus-driven outflows
@@
Wheeler et al. (2020) reported a collimated and extended molecular outflow in the E nucleus traced by CO transitions (J). The outflow found in the present work (SEO), launched from the E nucleus, is consistent with the direction of the outflow observed in the cold molecular gas, blueshifted toward the south direction. Wheeler et al. (2020) reported a mass for the blueshifted outflow of 2.7 M⊙, which, under the assumption that the CO outflow is associated with the SE outflow detected in our data, would lead to a hot-to-cold molecular gas ratio of . This is compatible with the value we inferred in Sect. 6.1.1 for the outflow launched by the W nucleus.
@@
We attributed the NW bubble to the E nucleus, since this structure could represent the counterpart of the blushifted outflow detected in the SE direction. However, unlike the HGO, in this case there is no clear blushifted and redshifted part of the outflow. This may be due to different gas ejection events, or to the fact that the redshifted part of the outflow is more obscured. Another possibility is that we included part of the redshifted outflow launched from the E nucleus in the HGO, since the two outflows may overlap, as we can see in Fig. 5c. From Fig. 19 we see that assigns a higher flux weight to the clouds of the HGO near the W nucleus. This suggests the possibility of contamination from the redshifted counterpart of the SEO.
@@
We find that the mass of the ionized gas in the NW bubble, assuming ne = 170 cm-3, is M⊙, with a total energy of erg. Lockhart et al. (2015) estimated the properties of the NW bubble using H+[NII] narrow band imaging obtained with HST/WC3 with a resolution of 0.0396”/pixel, corresponding to 15 pc. They used assumptions different from ours on the density, on velocity of the bubble and also on the geometry. In fact, they adopted ne = 10 cm-3, about one order of magnitude lower than that used in this work, a velocity of 250 km s-1, around a factor 2 lower and a shell-like geometry. Despite these different assumptions, their derived outflow mass ( M⊙) and energy ( erg) are in agreement, within a factor of two, with our results.
@@
We also compared the outflows detected with NIRSpec with the larger scale outflows observed with MUSE in Perna et al. (2020). In that case, the most extreme velocities for the ionized outflow are observed along PA = 138° 222Measured anti-clockwise from the north.; these extreme motions are aligned with X-ray emission and with the minor axis of the stellar disk. The SEO and the NWB are found along the same PA, but at ¡ 1 kpc scales, which might suggest that they are related to the kiloparsec-scale atomic outflow presented in Perna et al. (2020). The ionized, large-scale bubble observed in MUSE (kpc) and the NW bubble could be, also in this case, the results of different outflow episodes. These multiple bubbles at different distances are very similar to those predicted by the TNG50 cosmological simulation as a result of AGN feedback (Nelson et al. 2019) and to those observed in NGC 1275 (e.g., Fabian et al. 2003; Sanders & Fabian 2007) and in the Tea Cup (Venturi et al. 2023).
6.2 The launching mechanism